OpenStudy (anonymous):

I don't know if I have the info I need to get the number I am looking for. At work I have a report that shows the number of medications scanned. Total= 5000 meds scanned( this includes the whole hospital nurses and therapists) It also shows 750 of these were scanned outside of a time variance, which would equal 0.15% of the total. I count how many of these outside of the time variance were scanned by therapists alone. Say 100 of these were scanned by therapist alone. So my therapist % would equal 0.02% of the total. If 5000 is the total for the nurses and therapists, how can I figure out how

OpenStudy (dumbcow):

one thing i noticed 750/5000 = .15 = 15% not 0.15% so the 100 scanned by therapist would be 2% of total ok now im not sure what you need figured out?

OpenStudy (anonymous):

yes, dumbcow is right, And I can't see your whole question. What do you need figured out?

OpenStudy (anonymous):

Thanks for correcting me. Math is not my strong suit. If 5000 is the total of everyone, I want to figure out what the total for therapist alone are. I know there are 100 outside of the variance window, but I would like to know how many of that 5000 total are therapist? ( the 5000 is the total of all that are within and outside of the variance. So I would like to know the same for therapist alone if it can be derived from the info I have.) Thanks.

OpenStudy (dumbcow):

ok so if 750 are outside time variance, then 4250 are inside so the total is 5000 To know how many of those 4250 are therapists you have to know the ratio of nurses to therapists. For example how did you know 100 out of the 750 were therapists, if that represents a consistent ratio then num of therapists would be 4250*100/750 = 566.6. Also are therapists more likely to be inside or outside the time variance?

OpenStudy (anonymous):

if totally 750 were scanned outside the time variance, it means that 5000-750 = 4250 were scanned within the time variance. So, it 100 were scanned outside time variance by therapists, alone, then they scanned 100/750 th of 4250 within the time variance.

OpenStudy (anonymous):

I print the report and it has the med and the names of the person who scanned it. I go through the report manually and count how many were therapists out of all the ones listed on the report. No, therapist are not necessarily more likely to be outside of the variance.

OpenStudy (dumbcow):

ok then if you just count the ones outside time variance then you could estimate the therapists inside Therapists Out = Total Out *(Therapist In/Total In)

OpenStudy (dumbcow):

oops i got In and out mixed up Therapists In = Total In *(Therapist Out/Total Out) does that help

OpenStudy (anonymous):

this is starting to sound like a dirty movie :D

OpenStudy (anonymous):

Can you put those numbers into that equation for me. I trying get it. I so much appreciate your help. If I can figure that out it will save me so much time on another report I have to hand track.

OpenStudy (anonymous):

LOL, dhatraditya

OpenStudy (dumbcow):

Sure TOTAL = 5000 Total Out = 750 Total In = 5000 - 750 = 4250 Therapist Out = 100 Therapists In = 4250*(100/750) = 4250*(.1333) = 566.6 Of course this would be just an estimate so expect to be off by +- 20

OpenStudy (anonymous):

I was just starting to see it. That number sounds about right. You would think they could make a report for what we need but I guess it makes me use my brain. Thank you so much for your help. It will help me.

OpenStudy (dumbcow):

haha yeah, you could get someone to put that formula into excel or something for you anyway your welcome